If 3 balls are simultaneously removed from the urn, then:

I. Expect to extract 1.5 red balls.

II The probability of not extracting red balls is 0.0833

a. None is correct
b. Both are correct
c. Only II is correct
d. Only I is correct
Answer #1

Since the number of red balls is 50% of the total balls in the urn. So if 3 balls are drawn simultaneously, then we can expect to extract 1.5 red balls.

And the probability that no ball drawn is red = (5/10)(4/9)(3/8) = 1/12 = 0.083

Hence, both the given statements are correct.

Hence the correct answer is the second option b.